# NoiseSDF2NoiseSDF: Learning Clean Neural Fields from Noisy Supervision

### v1

@@ Abstract
Reconstructing accurate implicit surface representations from point clouds remains a challenging task, particularly when data is captured using low-quality scanning devices. These point clouds often contain substantial noise, leading to inaccurate surface reconstructions. Inspired by the Noise2Noise paradigm for 2D images, we introduce NoiseSDF2NoiseSDF, a novel method designed to extend this concept to 3D neural fields. Our approach enables learning clean neural SDFs directly from noisy point clouds through noisy supervision by minimizing the MSE loss between noisy SDF representations, allowing the network to implicitly denoise and refine surface estimations. We evaluate the effectiveness of NoiseSDF2NoiseSDF on benchmarks, including the ShapeNet, ABC, Famous, and Real datasets. Experimental results demonstrate that our framework significantly improves surface reconstruction quality from noisy inputs.
@@ 1 Introduction
Learning from imperfect targets [49, 19, 25, 5, 51] is a fundamental challenge in machine learning, particularly when obtaining clean training labels is impractical or unfeasible. In image processing, the pioneering work of Noise2Noise [25] demonstrated that image restoration could effectively be achieved by observing multiple corrupted instances of the same scene. Specifically, this method leverages the principle that pixel values at identical coordinates in different noisy images ideally represent the same underlying true signal. Consequently, the model learns to restore clean images by simply minimizing a straightforward MSE loss between noisy observations. See Figure 1 (a).
Extending Noise2Noise principles to 3D point clouds [20, 31], however, poses significant challenges due to their inherently unstructured nature. Unlike images organized on regular grids, point clouds exhibit deviations across all spatial coordinates without the benefit of a stable reference framework. This fundamental difference renders a direct extension of unsupervised image denoisers impractical. Standard loss functions such as Mean Squared Error (MSE) prove ineffective, necessitating specialized loss functions like Earth Mover’s Distance (EMD) to capture geometric correspondences and spatial distributions inherent in point cloud data, see Figure 1 (b).
@@
Recent advances in surface reconstruction have introduced neural fields, such as neural Signed Distance Function (neural SDF) [35, 32, 13, 52], which are capable of predicting continuous SDF values for any given 3D coordinate. Our key observation is that neural SDF, which encodes the SDF mapping 3D coordinates to scalar distance values for 3D shape, exhibits a conceptual parallel to the mapping between pixel coordinates and pixel intensities in 2D images, as shown in Figure 1 (c). Building on this analogy, we hypothesize that neural SDFs can be denoised by directly using noisy SDF observations with the same MSE loss strategy inspired by the Noise2Noise principle in image restoration.
In this work, we explore the use of noisy-target supervision in neural SDFs for surface reconstruction from noisy 3D point clouds. We propose NoiseSDF2NoiseSDF, an adaptation of the Noise2Noise framework applied to neural fields. The workflow of our NoiseSDF2NoiseSDF is illustrated in Figure 2. The network first takes independently corrupted point clouds as input to predict the underlying clean SDF values. Instead of using clean SDFs as ground truth, we employ another noisy neural SDF, which is generated by off-the-shelf point-to-SDF methods, as the supervision target. We then minimize the discrepancy between the predicted SDF output and the noisy SDF target using MSE loss. The network learns to suppress noise and improve consistency across SDF values, leading to clean neural representations.
To evaluate the effectiveness of NoiseSDF2NoiseSDF, we conduct comprehensive experiments across benchmark datasets, including ShapeNet [8], ABC [23], Famous [15], and Real [15]. Our experimental results demonstrate that neural SDFs can indeed be denoised effectively by employing mean squared error loss directly between their noisy representations. This finding confirms our central hypothesis: neural SDFs can learn to produce cleaner outputs simply by observing and minimizing discrepancies among noisy neural fields, effectively extending the Noise2Noise paradigm into the domain of 3D neural surface reconstruction. Our approach eliminates the need for clean training data, making it practical and scalable for real-world scenarios where acquiring perfect data is difficult or infeasible.
@@ 2 Related Work
Noise2Noise. The Noise2Noise (N2N) framework [25] has significantly influenced recent image denoising. By leveraging pairs of noisy observations of the same scene, the N2N framework learns to predict one noisy realization from another via pixel-wise correspondence. Subsequent methods like Noise2Void [24], Noise2Self [3] employ blind-spot masking techniques, training models directly on individual noisy images without pairs. Noise2Same [44] derives self-supervised loss bounds to eliminate the blind-spot restriction altogether. Self2Self [39] and Neighbor2Neighbor [21] exploit internal image redundancy, employing dropout or pixel resampling to train directly on single noisy observations without explicit noise modeling. Noisier2Noise [34] extends N2N to explicitly introduce additional synthetic noise, learning to map noisier images back to their original noisy versions.
Extending the Noise2Noise framework to 3D [20, 43, 31, 41] is challenging due to the unordered nature of point clouds. Methods like TotalDenoising [20] and Noise2Noise Mapping [31] address this by leveraging local geometric correspondences instead of exact point matches, replacing MSE with Earth Mover’s Distance (EMD) loss to better align noisy point clouds with the underlying surface. In our work, we exploit the structural similarities between neural fields and images, proposing a Noise2Noise denoising framework for 3D SDFs with MSE loss.
@@
Implicit Surface Reconstruction. Learning implicit surfaces from point clouds has seen significant advances. Overfitting-based methods optimize a neural implicit function for a single point cloud, often with ground-truth SDFs, normals or geometric constraints and physical priors. For example, SAL [1], SALD [2], and Sign-SAL [50] use point proximity and self-similarity cues. Gradient regularization techniques like IGR [18], DiGS [4], and Neural-Pull [30] improve stability and detail. Extensions such as SAP [37], LPI [9], and Implicit Filtering-Net [26] enhance reconstruction under sparse sampling and complex geometry. While accurate, these methods are typically sensitive to noise. Robust variants (e.g., SAP [37], PGR [27], Neural-IMLS [42], Noise2Noise Mapping [31] and LocalN2NM [10]) address this via smoothing, denoising priors, or self-supervision.
@@
In contrast to overfitting approaches, data-driven methods learn shape priors from large datasets. Global-latent methods, such as OCCNet [32], IM-NET [11], and DeepSDF [35], encode entire shapes into fixed-length latent codes, capturing overall semantics but often over-smoothing details. Local prior methods improve expressiveness by operating at finer scales. Grid-based approaches divide space into cells and learn small implicit functions per cell (ConvOccNet [36], SSRNet [33], Local Implicit Grid [17], Deep Local Shapes [7]). Patch-based methods segment point clouds into local regions and learn shared atomic representations (PatchNets [40], POCO [6], neighborhood-based [22]). Hybrid methods combine global context with local detail. For instance, IF-Nets [12] and SG-NN [14] fuse PointNet features with voxel hierarchies or contrastive scene priors. Recent transformer-based models (ShapeFormer [45], 3DILG [47], 3DS2V [48], LaGeM [46]) leverage self-attention for long-range structure modeling. Data-driven models trained with ground-truth SDFs are generally more robust to noise. Hybrid approaches like P2S [38] and PPSurf [16] use dual-branch networks to predict SDFs with explicit noise-level supervision. However, their performance degrades under extreme sparsity and noise of input point clouds. In contrast, our NoiseSDF2NoiseSDF framework embraces noise as a training signal, enabling reliable surface recovery from severely degraded inputs.
@@ 3 Preliminaries
Let $f_{\theta}$ denote a neural network parameterized by $\theta$, and let $x\in\mathbf{X}$ represent a spatial query coordinate. The network is designed to predict pixel intensities given a noisy image and the query coordinate. The prediction is written as:
@@
where $y_{i}$ is the noisy input image, $x$ is the queried pixel location, and $\hat{y}(x\mid y_{i})$ is the predicted pixel intensity at $x$.
The model is trained to minimize the expected squared error between the predicted pixel value and the corresponding pixel value in another independent noisy observation. The loss function is:
@@
where $y_{1},y_{2}$ are independent noisy observations of the same clean image, and $x\in\mathbf{X}$ is sampled uniformly from the image domain.
@@ 4 Method
Our proposed method investigates whether clean neural fields can be effectively learned by observing their noisy counterparts. Drawing inspiration from Noise2Noise, where noisy images directly serve as inputs and targets, we adapt this principle to learning neural fields from noisy point cloud data. In contrast to the direct usage of noisy images as input in traditional Noise2Noise setups, we employ a neural network conditioned on a noisy point cloud to predict neural SDFs at specific query coordinates. Rather than utilizing clean SDFs as supervision, our approach leverages noisy neural fields at identical coordinates derived from another independently noisy version of the same underlying shape. This ensures one-to-one correspondence between the predicted and target neural fields, allowing effective noise suppression through direct MSE loss minimization.
@@ 4.2 Implementation
Point Sampling. We first normalize the watertight meshes into a unit cube, then sample points from the surfaces to obtain the original point cloud $p$. Following the Noise2Noise Mapping protocol [25, 31], we apply zero-mean Gaussian noise to generate noisy point cloud pairs. The query point set consists of $50\%$ near-surface points and $50\%$ uniformly sampled points from the unit cube. To reduce dependency on the original clean surface, we directly use the two input noisy point clouds as the near-surface query points. Additionally, we uniformly sample $N$ points within the cube as spatial query points.
Denoising Network. Our SDF prediction network is built on 3DS2V [48]. Initially, a noisy point cloud $p_{1}$ is sampled and transformed into positional embeddings, which are then encoded into a set of latent codes through a cross-attention module. Subsequently, self-attention is applied to aggregate and exchange information across the latent set, enhancing feature integration. A cross-attention module then computes interpolation weights for the query point $q$. These interpolated feature vectors are processed through a fully connected layer to predict SDF values. The network is initialized using pretrained weights from 3DS2V. Without the denoising learning, the predicted SDF values are typically noisy due to the inherent noise in the input point clouds. During our denoising phase, the encoder remains fixed while only the decoder is optimized, improving both efficiency and convergence speed. Importantly, our contribution is not limited to the specific architecture of the SDF prediction network; other point-to-SDF methods could also be utilized.
Noisy Target. Given another paired noisy point cloud $p_{2}$, a Point2SDF method is required to predict noisy SDF values $s_{2}$ from it. In our implementation, we also employ a pretrained 3DS2V [48] model. This network accepts as input a noisy point cloud $p_{2}$ and a query point $q$, producing the corresponding noisy SDF scalar value at $q$. To ensure that all SDF targets are consistently noisy, we freeze its parameters during this process. Notably, this Point2SDF method is also not restricted to a specific model; any data-driven or overfitting Point2SDF methods could be seamlessly integrated. The Mean Squared Error (MSE) loss function is employed to minimize discrepancies between corresponding SDF values. This loss guides the optimization of the neural SDF network’s weights during the denoising training.
@@ 5 Experiment
We designed experiments to evaluate the performance of the NoiseSDF2NoiseSDF framework for surface reconstruction from raw noisy point clouds, assessing performance across different noise levels using a data-driven paradigm on large 3D shape datasets. We conducted ablation studies to validate key components and design choices.
@@ 5.1 Training Details
For optimization, we used the AdamW optimizer [29] with a fixed learning rate of $1\times 10^{-4}$. For resource usage, we trained on three Nvidia A100 GPUs with a batch size of 32 per GPU, taking approximately 15 hours for the ShapeNet dataset and 2.5 hours for the ABC dataset.
We sampled 2048 points from watertight meshes as the initial point cloud. Following the Noise2Noise Mapping [31], we applied Gaussian noise with standard deviations of $1\%$, $2\%$, online to generate noisy and sparse point cloud pairs. Additionally, we sampled 8192 query points online. The noise magnitude is defined with respect to both the point-cloud bounding-box size and the point density. For a fixed numeric noise level, a smaller bounding box amplifies the relative impact of the perturbation. All point clouds are normalized to the cubes $[-0.5,0.5]^{3}$ or $[-1,1]^{3}$. Furthermore, sparser point sets are more susceptible to noise. With only 2048 points, noise levels of $0.01$ and $0.02$ constitute severe perturbations irrespective of the bounding-box scale. The clean underlying surface is recovered from the denoised SDF using the Marching Cubes [28].
@@ 5.2 Datasets and Metrics
We trained our NoiseSDF2NoiseSDF network on the ShapeNet dataset following [48], using the same data split and preprocessing procedures. To evaluate denoising effectiveness and surface reconstruction quality, we used several metrics, including Intersection-over-Union (IoU), Chamfer Distance, F1 Score, and Normal Consistency.
IoU was computed based on occupancy predictions over densely sampled volumetric points. Following methods [26, 30], we sampled $1\times 10^{5}$ points from the reconstructed and ground-truth surfaces to compute the Chamfer Distance and F1 Score.
To further assess the generalization capability of NoiseSDF2NoiseSDF, we further trained our model on the ABC train set [23]—which was not used during the pretraining of 3DS2V—and then evaluated it on the ABC test set, as well as the Famous [15] and Real [15] datasets. We utilized the preprocessed datasets and data splits provided by [15, 16]. We reported evaluation metrics including Normal Consistency, Mesh Normal Consistency, Chamfer Distance, and F1 Score. All metrics reported above are evaluated on the reconstructed meshes. We excluded IoU from this benchmark because, under severe noise, many reconstructed meshes become non-watertight or heavily degenerated, making it infeasible to assign reliable inside/outside labels and rendering the IoU metric unreliable.

### v2

@@ Abstract
Reconstructing accurate implicit surface representations from point clouds remains a challenging task, particularly when data is captured using low-quality scanning devices. These point clouds often contain substantial noise, leading to inaccurate surface reconstructions. Inspired by the Noise2Noise paradigm for 2D images, we introduce NoiseSDF2NoiseSDF, a novel method designed to extend this concept to 3D neural fields. Our approach enables learning clean neural SDFs from noisy point clouds through noisy supervision by minimizing the MSE loss between noisy SDF representations, allowing the network to implicitly denoise and refine surface estimations. We evaluate the effectiveness of NoiseSDF2NoiseSDF on benchmarks, including the ShapeNet, ABC, Famous, and Real datasets. Experimental results demonstrate that our framework significantly improves surface reconstruction quality from noisy inputs.
@@ 1 Introduction
Learning from imperfect targets [51, 19, 27, 5, 54] is a fundamental challenge in machine learning, particularly when obtaining clean labels is impractical or unfeasible. In image processing, the pioneering work of Noise2Noise (N2N) [27] demonstrated that image restoration could effectively be achieved by observing multiple corrupted instances of the same image. Specifically, N2N leverages the principle that pixel values at identical coordinates in different noisy images ideally represent the same underlying true signal. Then, the model learns to restore clean images by minimizing a simple loss, such as mean squared error (MSE), between noisy observations, as shown in Figure 1 (a).
Extending N2N principles to 3D point clouds  [21, 33], however, poses inherent limitations due to their unstructured nature. Unlike images organized on regular grids, point clouds exhibit deviations across all spatial coordinates without the benefit of a stable reference framework. This fundamental difference renders a direct extension of N2N impractical. Standard loss functions such as MSE prove ineffective, resorting to specialized loss functions like Earth Mover’s Distance (EMD) to capture only soft geometric correspondences in point cloud data, see Figure 1 (b).
@@
Recent advances in 3D shape surface reconstruction have introduced neural fields, such as neural Signed Distance Function (SDF) [36, 34, 13, 55], which are capable of predicting continuous SDF values for any given query 3D coordinate. Our key observation is that neural SDF, which encodes the 3D shape as an SDF mapping from 3D coordinates to scalar distance values, exhibits a conceptual parallel to the mapping between pixel coordinates and pixel intensities in 2D images, as shown in Figure 1 (c). Building on this analogy, we hypothesize that neural SDFs can be denoised by using noisy SDF observations with the same MSE loss strategy inspired by the N2N principle in image restoration.
In contrast to paired noisy 2D images that can be readily acquired from cameras, 3D neural fields are typically learned through neural networks. Because image coordinates are discrete and finite, simply following the image-level N2N approach and denoising the SDF values only at specified coordinates, cannot ensure that the denoised SDF field remains complete and continuous over an infinite coordinate domain.
To address the above limitation, we propose training a neural field network with denoising capability to predict clean SDF values without being restricted to specific query coordinates. Specifically, we introduce NoiseSDF2NoiseSDF, a novel approach that employs noisy-target supervision within neural SDFs to enable robust surface reconstruction from noisy 3D point clouds. The denoising network takes independently corrupted point clouds as input and predicts the underlying clean SDF values. Instead of using clean SDFs as ground truth, we employ another noisy neural SDF, generated by off-the-shelf point-to-SDF methods, as the supervision target. We then minimize the discrepancy between the predicted SDF output and the noisy SDF target using MSE loss. Through this process, the network learns to suppress noise and improve consistency across SDF values, resulting in clean neural representations.
To evaluate the effectiveness of NoiseSDF2NoiseSDF, we conduct comprehensive experiments across benchmark datasets, including ShapeNet [8], ABC [25], Famous [15], and Real [15]. Our experimental results demonstrate that neural SDFs can indeed be denoised effectively by employing MSE loss directly between their noisy representations. This finding confirms our central hypothesis: neural SDFs can learn to produce cleaner outputs simply by observing and minimizing discrepancies among noisy neural fields, effectively extending the N2N paradigm into the domain of 3D shape surface reconstruction.
@@ 2 Related Work
Noise2Noise (N2N) [27] has significantly influenced recent image denoising. By leveraging pairs of noisy observations of the same scene, N2N learns to predict one noisy realization from another via pixel-wise correspondence. Subsequent methods like Noise2Void [26], Noise2Self [3] employ blind-spot masking techniques, training models directly on individual noisy images without pairs. Noise2Same [46] derives self-supervised loss bounds to eliminate the blind-spot restriction altogether. Self2Self [41] and Neighbor2Neighbor [23] exploit internal image redundancy, employing dropout or pixel resampling to train on single noisy observations without explicit noise modeling. Noisier2Noise [35] extends N2N to explicitly introduce additional synthetic noise, learning to map noisier images back to their original noisy versions.
Recent work has extended the N2N paradigm to the 3D domain, particularly to point clouds [21, 45, 33, 43]. TotalDenoising [21] and N2NM [33] employ soft local geometric correspondences with Earth Mover’s Distance (EMD) loss to align noisy point clouds with the underlying surface. However, these methods are unable to establish exact point matches.
To the best of our knowledge, no prior work has applied the N2N paradigm to the domain of 3D neural fields. We are the first to exploit the structural similarities between neural fields and images by proposing an N2N denoising framework for 3D SDFs using a simple MSE loss, which enables direct SDF matches.
Implicit Surface Reconstruction.
@@
Overfitting-based methods optimize a neural implicit function for a single shape through intensive test-time optimization. They often achieve high geometric fidelity on that specific object but lack generalization to new shapes. For example, SAL [1], SALD [2], and Sign-SAL [52] use point proximity and self-similarity cues. Gradient regularization techniques like IGR [18], DiGS [4], and Neural-Pull [32] improve stability and detail. Extensions such as SAP [39], LPI [9], and Implicit Filtering-Net [28] enhance reconstruction under sparse sampling and complex geometry. While accurate, these methods are typically sensitive to noise. Robust variants (e.g., SAP [39], PGR [29], Neural-IMLS [44], N2NM [33] and LocalN2NM [10]) address this via smoothing, denoising priors, or self-supervision.
@@
Data-driven methods learn from collections of shapes, allowing the model to infer implicit surfaces for previously unseen instances with efficient inference. For instance, global-latent methods, such as OCCNet [34], IM-NET [11], and DeepSDF [36], encode entire shapes into fixed-length global latent codes. Local-based methods improve expressiveness by operating at finer scales. Grid-based approaches divide space into cells and learn small implicit functions per cell (ConvOccNet [38], SSRNet [37], Local Implicit Grid [17], Deep Local Shapes [7]). Patch-based methods segment point clouds into local regions and learn shared atomic representations (PatchNets [42], POCO [6], neighborhood-based [24]). Hybrid methods combine global context with local detail. For instance, IF-Nets [12] and SG-NN [14] integrate local features within hierarchical representations. P2S [40] and PPSurf [16] use dual-branch networks to predict SDFs. Recent transformer-based models (ShapeFormer [47], 3DILG [49], 3DS2V [50], LaGeM [48]) leverage self-attention for long-range structure modeling. Since these methods are trained using ground-truth SDFs, their performance degrades when input point clouds are sparse or noisy. In this work, we demonstrate that clean neural fields can be learned under noisy supervision, enabling robust surface reconstruction from corrupted inputs.
@@ 3 Preliminaries
Let $f_{\theta}$ denote a denoising network parameterized by $\theta$, which was initialized using ImageNet pretraining [20], and let $x\in\mathbf{X}$ represent a spatial query coordinate. The network is designed to predict pixel intensities given a noisy image and the query coordinate. The prediction is written as:
@@
where $y_{i}$ is the noisy input image, $x$ is the queried pixel location, and $\hat{y}(x\mid y_{i})$ is the predicted pixel intensity at $x$. The model is trained to minimize the expected squared error between the predicted pixel value and the corresponding pixel value in another independent noisy observation. The loss function is:
@@
where $y_{1},y_{2}$ are independent noisy observations of the same clean image, and $x\in\mathbf{X}$ is sampled uniformly from the image domain. After training, given a noisy image, $f_{\theta}$ can predict a denoised version.
@@ 4 Method
Our proposed method investigates whether clean neural fields can be effectively learned by observing their noisy counterparts. Drawing inspiration from Noise2Noise, where noisy images directly serve as inputs and targets, we adapt this principle to learning neural fields from noisy point cloud data. In contrast to the direct usage of noisy images as input in traditional Noise2Noise setups, we employ a neural network conditioned on a noisy point cloud to predict neural SDFs at given query coordinates. Rather than utilizing clean SDFs as supervision, our approach leverages noisy neural fields at identical coordinates derived from another independently noisy version of the same underlying shape. This ensures one-to-one correspondence between the predicted and target neural fields, allowing effective noise suppression through direct MSE loss minimization.
@@ 4.2 Implementation
Point Sampling. We first normalize the watertight meshes into a unit cube, then sample points from the surfaces to obtain the original point cloud $p$. Following the Noise2Noise protocols [27, 33], we apply zero-mean Gaussian noise to generate noisy point cloud pairs. The query point set consists of $50\%$ near-surface points and $50\%$ uniformly sampled points from the unit cube. To reduce dependency on the original clean surface, we directly use the two input noisy point clouds as the near-surface query points. Additionally, we uniformly sample $N$ points within the cube as spatial query points.
Denoising Network. Our SDF prediction network is built on 3DS2V [50]. Initially, a noisy point cloud $p_{1}$ is sampled and transformed into positional embeddings, which are then encoded into a set of latent codes through a cross-attention module. Subsequently, self-attention is applied to aggregate and exchange information across the latent set, enhancing feature integration. A cross-attention module then computes interpolation weights for the query point $q$. These interpolated feature vectors are processed through a fully connected layer to predict SDF values. The network weights are initialized following Zhang et al. [50] to accelerate convergence.
Noisy Target. Given another paired noisy point cloud $p_{2}$, a Point2SDF method is required to predict noisy SDF values $s_{2}$ from it. In this implementation, we use the 3DS2V [50] with frozen parameters. Since it is not trained in a denoising setting, noisy inputs produce noisy SDF outputs. This network accepts as input a noisy point cloud $p_{2}$ and a query point $q$, producing the corresponding noisy SDF scalar value at $q$. To ensure that all SDF targets are consistently noisy, we freeze its parameters during this process.
Inference. After training, only the Denoising Network is used for inference. Given a noisy point cloud, the network predicts the denoised SDF values in a single forward pass. The clean underlying surface is then recovered from the denoised SDF using Marching Cubes [30].
@@ 5.1 Training Details
For optimization, we used the AdamW optimizer [31] with a fixed learning rate of $1\times 10^{-4}$. For resource usage, we trained on three Nvidia A100 GPUs with a batch size of 32 per GPU, taking approximately 15 hours for the ShapeNet dataset and 2.5 hours for the ABC dataset. We sampled 2048 points from watertight meshes as the initial point cloud. Following the N2NM [33], we applied Gaussian noise with standard deviations of $1\%$, $2\%$, online to generate noisy and sparse point cloud pairs. Additionally, we sampled 8192 query points online. The noise magnitude is defined with respect to both the point‑cloud bounding‑box size and the point density. For a fixed numeric noise level, a smaller bounding box amplifies the relative impact of the perturbation. All point clouds are normalized to the cubes $0.5,0.5]^{3}$ or $1,1]^{3}$. Furthermore, sparser point sets are more susceptible to noise. With only 2048 points, noise levels of $0.01$ and $0.02$ constitute severe perturbations irrespective of the bounding‑box scale.
@@ 5.2 Datasets and Metrics
We evaluate our NoiseSDF2NoiseSDF on ShapeNet following Zhang et al. [50]. To assess denoising effectiveness and surface reconstruction quality, we use metrics, including Intersection-over-Union (IoU), Chamfer Distance, F1 Score, and Normal Consistency (NC). IoU was computed based on occupancy predictions over densely sampled volumetric points. Following methods [28, 32], we sampled $1\times 10^{5}$ points from the reconstructed and ground-truth surfaces to compute the Chamfer Distance and F1 Score.
To further evaluate the generalization ability of our approach, we trained the model on the ABC training set [25] and tested it on the ABC test set, as well as the Famous [15] and Real [15] datasets. Importantly, neither of these datasets was used during 3DS2V’s training, so they can be regarded as out-of-distribution. Our model is trained solely with noisy targets generated by 3DS2V, without relying on any clean ground-truth from these datasets. We utilized the preprocessed datasets and data splits provided by Erler et al. [15, 16]. We reported evaluation metrics including Normal Consistency, Mesh Normal Consistency, Chamfer Distance, and F1 Score. All metrics reported above are evaluated on the reconstructed meshes. We excluded IoU from this benchmark because, under severe noise, many reconstructed meshes become non-watertight or heavily degenerated, making it infeasible to assign reliable inside/outside labels and rendering the IoU metric unreliable.